OpenStudy (anonymous):

you invest an amount of money in a mutual fund. after four years, the balance of your account is $1669.34. assuming the interest rate is 7.5% and compounded annually, what is the initial amount you invested?

jimthompson5910 (jim_thompson5910):

A = P(1+r/n)^(n*t) 1669.34 = x(1+0.075/1)^(1*4) solve for x

OpenStudy (anonymous):

is it 1250?

jimthompson5910 (jim_thompson5910):

yep, you got it

OpenStudy (anonymous):

thanks!